Grade 11Physical Chemistry

10 ml of 0.1N HClHCl is added to 990 ml solution of NaCl.The pH of resulting solution is

Shashank Raut
10 Years agoGrade 11

1 Answer

Ashutosh Mohan Sharma
10 Years ago
moles of H+ = 10* (0.1/1000)
= 10^-3
volume of solution = 10+ 990 = 1000

conc of H+= moles of H+/Total volume of solution = 10^-3/1000 = 10^-6
pH =-log(H+) = -log(10^-6) =6

note: presence of NaCl does not change pH of soolution as pH is dependent only on concentration of proton.